OpenStudy (anonymous):

Check this? Find the solution of the linear equation. 2/3 (6 - 3x) = (1 - x) + (3 - x) A. x = -8 B. x = 8 C. infinitely many solutions D. no solution <---

hero (hero):

@CrazyCountryGirl, can you show the work you did to arrive at no solution?

hero (hero):

Have you tried solving for x yet?

OpenStudy (anonymous):

They left before they could show me how.

hero (hero):

The best approach for this is to work with one side at a time. What is (1 - x) + (3 - x)?

hero (hero):

Keep in mind, you can group like terms together.

OpenStudy (anonymous):

Nevermind. @geerky42 said it was correct. :)

hero (hero):

Your answer choice is not correct.

geerky42 (geerky42):

We already discussed here: http://openstudy.com/study#/updates/5393a469e4b01c619df0ed34

hero (hero):

Discussing it doesn't mean the correct answer was found. It just means that it was discussed.

hero (hero):

Many mathematicians and scientists discuss things and sometimes they come up with a solution to a problem and sometimes they don't.

hero (hero):

Okay, @CountryGirl, would you like to add (1 - x) + (3 - x)?

geerky42 (geerky42):

No need to do, I made small mistake as we process on other question. Case is closed now

OpenStudy (anonymous):

What is x? How do I solve it? o_o

hero (hero):

@CrazyCountryGirl, you have to solve it one step at a time. I'm trying to walk you through the process.

geerky42 (geerky42):

I did said D, but turn out I add wrongly because we take slow step to step, so turn out to be C.

geerky42 (geerky42):

we already walked through, step to step. now we have to do again?

hero (hero):

So @geerky42, you're just going to give the answer like that? I don't care what you did last time. Whatever you did @CrazyCountryGirl didn't understand it. If she did, she wouldn't have posted the question a second time.

OpenStudy (anonymous):

I understood enough od it to know it wasn't A or B.

geerky42 (geerky42):

Look at link I gave, you will see that I did take step to step, we are doing that for almost two hour in fact, and at each step, she seem to understand. I don't give her answer instantly in other question

geerky42 (geerky42):

Want to go through for two hours, be my guest, at least she is getting some helps.

hero (hero):

@geerky42, I was helping her and you gave the answer. Which is very rude on your part. I could have helped her in ten minutes had you not gave the answer.

geerky42 (geerky42):

Like I said, we took step to step in other question. did you read much?

hero (hero):

Giving the answer doesn't help anyone learn anything. You tried to help her in the other question, she didn't get it. She posted the question again, then tagged me. In other words, give someone else a chance to help her and move on.

geerky42 (geerky42):

you jumped to conclusion and assume I gave her answer as soon as I saw her question... Look at other question via link I gave above here...

OpenStudy (anonymous):

I did understand some of it. The whole -x-x thing confused me.

hero (hero):

I was going to help you through it.

hero (hero):

Now, what I'm going to do is just post the steps I took to solve it.

OpenStudy (anonymous):

ok. :p

hero (hero):

2/3 (6 - 3x) = (1 - x) + (3 - x) First, I re-grouped the terms on the right side: 2/3 (6 - 3x) = (1 + 3) + (-x - x) 2/3 (6 - 3x) = 4 + (-2x) 2/3 (6 - 3x) = 4 - 2x Next I factored out 3 from (6 - 3x): 2(3)/3 (2 - x) = 4 - 2x Which simplifes to 2(2 - x) = 4 - 2x Then you divide both sides by 2: 2 - x = 2 - x At this point you realize that both sides are the same which means no matter what x value you input, the output will be the same on both sides. If you have any questions about any of the steps, please let me know.

OpenStudy (anonymous):

I get it! I have two more questions to open in another thread. Can you help me? :)